Grade 12th passAlgebra

Simplify and solve the following linear equations.
3(t – 3) = 5(2t + 1)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

3(t–3) = 5(2t + 1)
⇒ 3t- 9 = 10t + 5
⇒ 3t- 10t = 5 + 9
⇒-7t = 14
⇒t = 14/-7
⇒t = -2

Thanks